Please explain to me the counting principle in probability,what does factorials mean and why do we sometimes divide factorials?Please explain all the rules of counting principles with simple examples and when must I use which rule please?
1
Expert's answer
2019-11-07T11:41:25-0500


Fundamental counting Principle :


Let E1 and E2 are two events. The first event E1 happens in "m" different ways,. After the happening of E1, the event E2 happens in n different ways.


The number of ways that the two events can happens is m "\times" n


This principle can be extended to three or more events.


Factorial means:

Factorial means the product of an integer and all integers below that required number.

Example:

"4! = 4 \times 3 \times 2 \times 1"

we sometimes divide factorials: When we are trying to find the combinations or permutations, we can use to divide the factorial.


Example:

"\binom {n} {r } = \frac {n!}{(n-r)! r!}"


Examples for counting Principle:


(1). If we have 3 shirts and 4 pants, the number of ways wearing these shirt and pant combination is 4 x 3 = 12 ways


(2). There is three cities A, B and C, number of routs from city A to city B is 3 and city B to C is 4 ways.


Total number of ways to go from city A to city B = 3 x 4 = 12


Fundamental Rules


we can use the fundamental rule of multiplication, if we have independent events.


We can use the fundamental rule of addition, if we have dependent events.